# OLMES: A Standard for Language Model Evaluations

### v1

@@ 6 Discussion
In creating this evaluation standard, OLMES, we build on top of the various previous efforts on language model evaluation in the community – including previous work on language model evaluation standardization, the many open research reports disclosing how evaluation on LMs have been done, and the datasets that made OLMES possible, which we explicitly cite and acknowledge in our paper.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] In Discussion section.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [N/A]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] URL at the end of Introduction section.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] In Section 2.
    2. (b)
      Did you mention the license of the assets? [N/A] All models and datasets used can be research for research purposes.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] In Appendix F.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A OLMES: implementation details
  • For MMLU: use macro average (over 57 tasks) rather than micro average (over 14042 instances), following AI@Meta, (2024). This better represents the diversity of fields in the dataset, although in practice it does not generally make a big difference (see Figure 7).
  • When a model requires it, make sure to add the appropriate <bos> token at start of prompt (e.g., Gemma (Gemma Team et al.,, 2024)).
  • Restrict all inputs (with completions) to 2048 tokens for consistency across models [^9] For current tasks this is only exhausted for a few MMLU instances.
  • Use the default model precision when evaluating (i.e., avoid options like load_in_8bit unless it produces identical results)
  • Other than the original instruction line for MMLU Hendrycks et al., (2021), we do not add any extra instructions. This is in view of previous work finding the subject information from instructions makes little changes to model ranking (Alzahrani et al.,, 2024), and to reduce additional sources of variation in the prompt.
Note that computational details, like batch size and type/state of GPU, can affect floating point operations such that answer choice decisions can flip if they are very close. This is hard to avoid unless consider “ties” when answers are sufficiently close in confidence.
@@ Appendix B Detailed CF and MCF task scores
Tables 5 and  6 present detailed scores across all tasks, with both MCF and CF results (using the OLMES recommendations for CF normalization).
@@ C.1 Task formulation details
LLM evaluations started out using the CF approach for many tasks (Brown et al.,, 2020; Du et al.,, 2022; Smith et al.,, 2022; Chowdhery et al.,, 2022; Lieber et al.,, 2021), which is a more reasonable option for weaker models that struggle with the more natural MCF (Khatun and Brown,, 2024). The task formulation only very recently and gradually switched to the MCF approach when it became clear that the model could utilize it, producing higher scores (Robinson et al.,, 2023; OpenAI,, 2024; AI@Meta,, 2024).
The HELM study (Liang et al.,, 2023) included comparisons between the MCF (“joint”) and CF (“separate”) approaches, finding that certain models can really benefit from the MCF approach, although among the models in the original study it was really only the Anthropic-LM v4-s3 (52B) model which could take full advantage of it.
@@ C.2 CF normalization details
Tables  9,  10 and  11 show detailed comparisons of CF normalization on different models, for the various tasks.
@@
When evaluating the GPT-3 model(Brown et al.,, 2020), they worked around this issue by normalizing the log-probability by the number of tokens in the answer (similar to how loss is computed during training). They also noted that for a few datasets, it worked markedly better to instead “normalize” by dividing by LLM probability of the same answer string without the presence of the question (usually by just having a generic prefix like "Answer: <answer_string>"). This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al.,, 2021).
The Eleuther LM Evaluation Harness (Gao et al.,, 2023; Biderman et al.,, 2024) and some subsequent evaluations (e.g., the Llama models (Touvron et al., 2023a, )) have also used “per answer character” normalization, using the argumentation (Gao,, 2021; Biderman et al.,, 2024), that normalizing per token is problematic since it depends on the tokenizer. Since the purpose of the normalization is simply to rank the answer choices within themselves (keeping model and tokenizer fixed), this does not seem like a relevant argument, and indeed a normalization which changes the probability of “yes” vs “no” simply because the “no” token has fewer characters seem problematic. In practice, for tasks where answers either relatively long or similar in length, there are minor differences between these two length normalization.
The HELM study (Liang et al.,, 2023) included comparisons between these normalization approaches for a number of tasks and models (using the terms “separate” and “separate calibrated” for “token” and “pmi” respectively), eventually settling on a default choice for each, not unlike the choices in the GPT-3 report (Brown et al.,, 2020). The Eleuther LM Evaluation Harness generally reports two metrics for each multiple-choice task: acc (using the “none” normalization) and acc_norm (using the “character” normalization).
@@
OLMo-1.7-7B 45.5 49.3 48.5 51.2 pmi 79.2 81.5 79.7 71.1 char 85.9 83.8 85.9 85.6 none
@@
Table 10: Comparing CF normalization schemes (part 2)
@@
OLMo-1.7-7B 71.6 63.5 59.0 70.4 none 61.4 80.1 77.7 65.2 char 39.9 42.4 42.2 41.8 char
@@
Table 11: Comparing CF normalization schemes (part 3).
@@ C.2.2 Hybrid formulation
In CF, overall probability score could be quite misleading since it may heavily favor shorter answers with fewer tokens. Note that this would be different if the answer choices are actually listed before scoring the answer string, then most tokens (after the choice has been disambiguated by the first few tokens) would have probability near one. This “hybrid” formulation has been used in some cases, but usually scores in between the CF and MCF approaches (Wiegreffe et al.,, 2023). However, this hybrid approach is not popular in evaluation standardization efforts like the Open LLM Leaderboard, HELM, or when used to evaluate models during development, so it is not a focus in OLMES.

### v2

@@ Acknowledgments
In creating this evaluation standard, OLMES, we build on top of the various previous efforts on language model evaluation in the community – including previous work on language model evaluation standardization, the many open research reports disclosing how evaluation on LLMs have been done, and the datasets that made OLMES possible, which we explicitly cite and acknowledge in our paper.
Limitations
The current version of OLMES is focused on providing guidance useful for LLM evaluation during the training stage and for comparing final base models, which provides important insights into the potential of such models before further tuning (e.g., instruction-tuning, safety-tuning). Interesting directions for future work include looking into evaluations targeted at accessing the effectiveness of various kinds of model tuning, as well as evaluation for multi-modal models.
This paper focuses on design choices in evaluating language models with multiple-choice tasks. While the suite of multiple-choice tasks used in this work includes questions on science, various types of commonsense, factual knowledge, and covers a range of topics (MMLU alone covers 57 subjects), of varying difficulty, an important future direction would be to apply the same principles in OLMES (e.g., prompt formatting, curated few-shot examples) to generative tasks and chain-of-thought prompting.
While the recommendations in OLMES are well-considered, justified and practical, they do not cover all plausible variants of presenting a task. See Appendix A for further discussion, showing how performance measured using OLMES is stable and consistent when subject to small changes in prompt wording or the selection of few-shot examples, within the general recommendations. Larger differences would be expected when diverging from OLMES recommendations such as by using unnatural prompts e.g., using rare symbols as answer labels, or randomly sampled few-shot examples which could run into skewed label distribution covered in few-shot examples or include noisy examples from train sets. We leave evaluating the robustness of models under adversarial setups as a topic for future work.
Ethical considerations
This study involves the use of large-scale language models. We only use their outputs to obtain their answers to questions in commonly used multiple-choice datasets, therefore we do not foresee any ethical issues with their use for the research presented in this work.
Appendix A FAQs
  • Q:
    Are there existing established protocols that differ from the proposed ones? Will they cause resistance from the research community in adopting the OLMES standard?
  • The lack of “established protocols” is precisely the issue that our work addresses. As discussed in our Introduction and Related work sections, existing efforts either present different variations as open choices to users or are underspecified or under-documented (e.g., the rationale behind default choice is unexplained) and thus not followed by others in subsequent work. Therefore, there is no existing established protocol in the community – researchers across the field use different evaluation setups, with reasons behind their choices left unexplained, leading to different results and conclusions. We illustrate this in Table 1, with Table 14 showing an extended version to include score variations across different references on OpenbookQA in addition to ARC-Challenge.
    This is the main motivation for the standard, to reconcile the differences in practices so that scores reported in papers can be meaningfully interpreted and compared (as we show, a statement like “score on 25-shot ARC Challenge” is woefully underspecified, whereas “score on ARC Challenge using OLMES” is a well-defined number without any ambiguity). We also justify each decision we make so that the community can, for the first time, appreciate the rationale behind the setups and thus encourage broad adoption.
  • Q:
    What is novel about OLMES?
  • Building on the many existing works that introduce new methodologies (e.g., new way of prompting, probability normalization, etc), OLMES is the first work of its kind to provide a completely open, practical, reproducible, and documented evaluation standard with justified choices so that results across research work can be meaningfully compared. This fills an important gap in current research on LLMs – the adoption of OLMES by model developers and other researchers will help unify evaluation practices in the field for the first time, significantly shifting current research paradigms.
  • Q:
    Why is OLMES more principled than trying a range of settings?
  • Rather than having to train a model from scratch to discover patterns in task formulation, run the different settings to choose a normalization scheme, or delve into the same literature again to study the variants, the community can now directly build upon the various choices in OLMES.
    We hope to guide the community towards more well-documented and justifiable chosen evaluation settings like OLMES without having to go through trying a mix of less informed choices (which we argue should be avoided altogether). Through extensive literature review and experimentation, we observe that some settings provide better signals than others, and document them in this work to guide the community to use them, a few examples include:
    • CF gives a clearer signal early in training, which is helpful for developers to cheaply make modeling decisions. On the other hand, MCF is a better indicator of performance later on. CF often works better for weaker models while MCF is at random, and MCF is a better representation of task performance for stronger models.
    • Few-shot prompting is an effective and universal way to convey a task to an LLM (more stable learning curve than 0-shot) but going beyond 5 shots generally does not provide meaningful differences in scores.
    • For probability normalization, in BoolQ the only answer choices are “yes” or “no” which are single tokens, therefore no length normalization is needed. Even if for some models, the “character” normalization has slightly better performance on BoolQ (see Table 9), one should note that this is an accidental side effect of “yes” having one more character than “no” and indeed a normalization which changes the probability of “yes” vs “no” simply because the “no” token has fewer characters seem problematic.
    Not only are recommendations in OLMES backed by both existing literature and new experimental results, there is also little difference between the OLMES recommendation and the empirically best (“oracle”) normalization for each task and model, see “diff oracle” column. We argue that adopting such an approach is a better practice than blindly optimizing for the best performance e.g., problematically using “character” normalization for BoolQ.
  • Q:
    Including a broader range of datasets?
  • The focus on multiple-choice datasets in OLMES is motivated by their frequent use in evaluating base LLMs, where the evaluation seems straightforward (did the model predict the right answer?) but in practice, a statement like “model X scores Y on ARC Challenge” is generally uninterpretable (with unspecified details and cannot be meaningfully compared across references) without a clear evaluation standard like OLMES. In this paper, we focused on datasets chosen to provide guidance useful for the training stage and evaluation of base models, which provides important insights into the potential of such models before further tuning (e.g., instruction-tuning).
    Note that the fundamental principles of OLMES as introduced, generalize to any dataset of interest. Rather than viewing what we have illustrated in our paper as a fixed set, our goal is to use that as an illustration and empower researchers to move towards reproducible evaluation by applying OLMES to any dataset of interest suited for their own work.
  • Q:
    Evaluating on more models? What are some valuable insights from extended experiments?
  • We provide additional results in Appendix D, Table 13 with additional models. Evaluating different models using OLMES provides valuable insights for understanding LLMs and model development. For instance, within each batch of model release by developers, models of bigger size perform better than smaller ones (see average scores of Pythia-6.7B outperforms Pythia-1B, OLMo-7B outperforms OLMo-1B, Llama2-13B outperforms Llama2-7B, Gemma2-9B outperforms Gemma2-2B). However, size is not the only way to get to a stronger model, evaluating on OLMES also allow the community meaningful comparison of models to witness the effect of model improvement via better training data, model architecture, as well as other improved approaches as researchers iterate on their models e.g., OLMo-7B-0424’s improvement over initial OLMo-7B; Llama3-8B’s improvement over Llama2-7B and even Llama2-13B, Mixtral-8x7B-v0.1 outperforming the Mistral model, aligning with the insights reported in these model releases documenting their improved recipes and innovations for better models. Further, OLMES also gives meaningful comparison of model performance as researchers experiment to reduce computational costs, e.g., our results align with the original DeepSeekMoE paper where they “scale up DeepSeekMoE to 16B parameters and show that it achieves comparable performance with LLaMA2 7B, with only about 40% of computations”. All these underscore the applicability and value of OLMES in supporting unified evaluation as the field progress towards better models, as an open, well-documented, practical and reproducible evaluation standard. We make all prompts, examples, and code used for OLMES openly available, and encourage researchers to try it for any model of their interest be it one they are studying or building.
  • Q:
    How does OLMES stay relevant in the rapid evolution of AI and LLMs?
  • We have been continuously looking out for new LLMs and evaluating them using OLMES, showing that the same guiding principles still apply as best practices providing a systematic, comparable approach. See extended evaluation results in Table 13.
    As the field moves forward, we look forward to applying the principles of OLMES to more benchmarks and evaluating newer models using OLMES. While we are working on extending OLMES, we do not anticipate revisions to the currently established recommendations in OLMES any time soon as the guiding principles are built on top of a rich literature of existing work over the years and will likely remain relevant in the community for a while in the near future.
  • Q:
    Why not use prompting techniques such as CoT or self-reflection?
  • While these prompting strategies have shown to be useful for instruction-tuned models, they tend to be much less effective on base models, which is a focus of this work. E.g., some experiments we performed with MMLU showed that various CoT prompts (both zero-shot and few-shot) have a positive boost on instruction-tuned models (like Llama-3.1-8B-Instruct), but tend to lower the scores a bit for base models (like Llama-3.1-8B).
  • Q:
    How do you ensure that formatting settings are fair to all models?
  • Supported by reviewing common evaluation practices in the community and empirical evidence across a wide range of models, the recommendations we make are at least as reasonable and fair as the myriad of settings that have been used in the literature.
    If a model is peculiar in any specific way (e.g., only able to do multiple choice questions with one type of answer label like “1.” or “2.”), it is not the goal of OLMES to tailor to such peculiarities as this standard is intended to be applied across a range of models and to encourage the development of models that produce reasonable outputs given any reasonable input.
  • Q:
    What happens when there are minor variants to OLMES?
  • Through OLMES, we provide best practices to evaluate language models and justify our choices. Our choices are mostly aligned with common practices in LLM evaluations, but with defining standards in formatting, choice of in-context examples, probability normalizations, and task formulation. In the process, we accounted for many factors, taking into consideration the robustness of OLMES under minor variations. We discuss some of these considerations here.
    [Part 1] Order of presenting the options A/B/C/D:
    The order of presenting the multiple-choice options A/B/C/D does not apply to CF since each answer is processed independently. For MCF it is indeed a confounder that some (especially weaker) models might highly prefer a given label (like B). The benchmarks in OLMES are generally balanced such that such a model would not be much better than random. Further, if this happens, CF would generally get a better score in such cases and OLMES would use that score in its final output. Therefore, having a setting where we use both CF (not affected by the order of options) and MCF (where the order of options may matter) makes sure the final metric will not be hugely affected by such factors. We considered applying more rigorous measures (like running all cyclic permutations of answer choices) but decided for practical reasons, the extra processing time and complexity were not worth the minor improvements in robustness (as one consideration of OLMES is also to be a practical standard that does not take unnecessarily more compute than is needed).
    [Part 2] Minor variations in prompt wording or few-shot examples:
    To address potential concerns on minor variations in prompt wording or few-shot examples, we evaluated under three additional settings, while adhering to the general principles in OLMES:
    Variant 1 (minor variation in prompting):
    3 changes to OLMES prompt format - (1) change the label and text separator from “.” to “)”, (2) insert an additional new line before the answer descriptor, (3) change the “Answer” descriptor to “Correct answer”
    Variant 2 (varying few-shot examples):
    Create a different set of curated few-shot examples by changing 3 out of the 5 in-context examples to new ones that are different from those in OLMES. In picking the new few-shot examples, the same recommendations were followed to ensure diversity in the examples and that they cover the label space.
    Variant 3 (minor variation in prompting + varying few-shot examples):
    Apply changes in both Variants 1 and 2 together.
    We report these additional results in Table 5. Following EleutherAI in calculating standard error (https://github.com/EleutherAI/lm-evaluation-harness/blob/ebe7226ebfb8d11a9fb8d6b53eb65891f895c633/lm_eval/api/metrics.py#L288), in the additional results, we also incorporated bounds on standard error in our evaluations using OLMES (see “std err” column). This provides a statistical bound on the degree of variation in reported numbers and illustrates that while any performance metric should be interpreted to have slight variants (e.g., < 2.5%), the scenario where a model underperforms significantly due to minor variants is unlikely statistically.
    model ARC_E orig var 1 var 2 var 3 avg diff std err
    Pythia-1B 63.4 63.3 62.7 62.7 63.0 0.4 1.5
    Llama2-7B 84.0 84.4 83.4 84.4 84.0 0.0 1.2
    DeepSeek-7B 80.6 80.9 80.5 80.4 80.6 0.0 1.3
    Gemma2-2B $84.3^{\dagger}$ $83.2^{\dagger}$ $83.9^{\dagger}$ $82.8^{\dagger}$ 83.5 0.8 1.2
    Llama3-8B $92.4^{\dagger}$ $92.5^{\dagger}$ $92.3^{\dagger}$ $93.1^{\dagger}$ 92.6 0.2 0.8
    model OBQA orig var 1 var 2 var 3 avg diff std err
    Pythia-1B 40.4 38.6 39.4 37.6 39.0 1.4 2.2
    Llama2-7B 57.8 57.2 55.2 57.4 56.9 0.9 2.2
    DeepSeek-7B $62.2^{\dagger}$ $61.0^{\dagger}$ $61.6^{\dagger}$ $63.2^{\dagger}$ 62.0 0.2 2.2
    Gemma2-2B $68.8^{\dagger}$ $67.2^{\dagger}$ $68.8^{\dagger}$ $67.0^{\dagger}$ 68.0 0.8 2.1
    Llama3-8B $77.2^{\dagger}$ $76.8^{\dagger}$ $78.8^{\dagger}$ $77.8^{\dagger}$ 77.7 0.5 1.9
    model PIQA orig var 1 var 2 var 3 avg diff std err
    Pythia-1B 68.9 69.2 69.2 69.3 69.1 0.2 1.5
    Llama2-7B 77.5 77.2 77.7 77.8 77.5 0.0 1.3
    DeepSeek-7B 79.3 78.8 80.9 80.6 79.9 0.6 1.3
    Gemma2-2B 78.5 77.8 79.3 78.5 78.5 0.0 1.3
    Llama3-8B 81.6 80.7 82.4 82.8 81.9 0.3 1.2
    Table 5: Extended results comparing using OLMES (orig) and when the setting is subjected to minor variations in prompt wording (var1), few-shot examples (var2), or both (var3). $\dagger$ indicates the use of MCF. The “avg” score obtained via averaging orig, var1, var2, and var3 results is often within 1% of that obtained by the original OLMES setup (orig). We report the observed differences between averaging the 4 setups (“avg”) and directly using OLMES (orig) in the “diff” column, illustrating the minor differences (often <1%) do not justify the 4 times more compute needed, against the “practical” consideration in OLMES.
    The additional results show that differences in performance between averaging variations vs. using the OLMES setup directly were generally minimal, typically less than 1 percent (the largest difference seen is 1.4%). This suggests that performance measured using OLMES is quite stable and consistent when subject to small changes in prompt wording or the selection of few-shot examples, within the general recommendations. Note that these variants still format the instances in natural ways and are slight modifications of the original settings of OLMES, still adhering to the general principles such as instance formatting that clarifies the task in a natural way and choice of in-context examples to cover a range of examples and different answer labels. Larger differences would be expected when diverging from OLMES recommendations such as by using unnatural prompts e.g., using rare symbols as answer labels, or randomly sampled few-shot examples which could run into skewed label distribution covered in few-shot examples or include noisy examples from train sets. We observe that current successful language models are generally robust to the OLMES evaluation standard. OLMES has been informed by prior efforts like HELM and Eleuther LM Evaluation Harness, therefore the prompts are designed to be natural, and suitable for evaluating language models.
@@ Appendix B Detailed CF and MCF task scores
Tables 6 and  7 present detailed scores across all tasks, with both MCF and CF results (using the OLMES recommendations for CF normalization).
@@ C.1 Task formulation details
LLM evaluations started out using the CF approach for many tasks (Brown et al., 2020; Du et al., 2022; Smith et al., 2022; Chowdhery et al., 2022; Lieber et al., 2021), which is a more reasonable option for weaker models that struggle with the more natural MCF (Khatun and Brown, 2024). The task formulation only very recently and gradually switched to the MCF approach when it became clear that the model could utilize it, producing higher scores (Robinson et al., 2023; OpenAI, 2024; AI@Meta, 2024).
The HELM study (Liang et al., 2023) included comparisons between the MCF (“joint”) and CF (“separate”) approaches, finding that certain models can really benefit from the MCF approach, although among the models in the original study it was really only the Anthropic-LM v4-s3 (52B) model which could take full advantage of it.
@@ C.2 CF normalization details
Tables  10,  11 and  12 show detailed comparisons of CF normalization on different models, for the various tasks.
@@
When evaluating the GPT-3 model (Brown et al., 2020), they worked around this issue by normalizing the log-probability by the number of tokens in the answer (similar to how loss is computed during training). They also noted that for a few datasets, it worked markedly better to instead “normalize” by dividing by LLM probability of the same answer string without the presence of the question (usually by just having a generic prefix like "Answer: <answer_string>"). This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al., 2021).
The Eleuther LM Evaluation Harness (Gao et al., 2023; Biderman et al., 2024) and some subsequent evaluations (e.g., the Llama models (Touvron et al., 2023a)) have also used “per answer character” normalization, using the argumentation (Gao, 2021; Biderman et al., 2024), that normalizing per token is problematic since it depends on the tokenizer. Since the purpose of the normalization is simply to rank the answer choices within themselves (keeping model and tokenizer fixed), this does not seem like a relevant argument, and indeed a normalization which changes the probability of “yes” vs “no” simply because the “no” token has fewer characters seem problematic. In practice, for tasks where answers are either relatively long or similar in length, there are minor differences between these two length normalizations.
The HELM study (Liang et al., 2023) included comparisons between these normalization approaches for a number of tasks and models (using the terms “separate” and “separate calibrated” for “token” and “pmi” respectively), eventually settling on a default choice for each, not unlike the choices in the GPT-3 report (Brown et al., 2020). The Eleuther LM Evaluation Harness generally reports two metrics for each multiple-choice task: acc (using the “none” normalization) and acc_norm (using the “character” normalization).
@@
OLMo-7B-0424 45.5 49.3 48.5 51.2 pmi 79.2 81.5 79.7 71.1 char 85.9 83.8 85.9 85.6 none
@@
Table 10: Comparing CF normalization schemes (part 1).
@@
OLMo-7B-0424 71.6 63.5 59.0 70.4 none 61.4 80.1 77.7 65.2 char 39.9 42.4 42.2 41.8 char
@@
Table 11: Comparing CF normalization schemes (part 2)
@@ C.2.2 Hybrid formulation
In CF, overall probability score could be quite misleading since it may heavily favor shorter answers with fewer tokens. Note that this would be different if the answer choices are actually listed before scoring the answer string, then most tokens (after the choice has been disambiguated by the first few tokens) would have probability near one. This “hybrid” formulation has been used in some cases, but usually scores in between the CF and MCF approaches (Wiegreffe et al., 2023). However, this hybrid approach is not popular in evaluation standardization efforts like the Open LLM Leaderboard, HELM, or when used to evaluate models during development, so it is not a focus in OLMES.